Find the basies of the column space and null space of the matrix.
A =

1 −2 5 −3 4
2 3 1 −4 2
.
• Find the the rank(A) and Nullity(A)
• If the matrix is augmented by a column what would be the rank and nullity of the
matrix?
1
Expert's answer
2021-09-20T10:55:11-0400

rref(A)= |1 0 17/7 -17/7 16/7|

| 0 1 -9/7 2/7 -6/7|

Basis of column space are 1st and second column

Column space=

"\begin{bmatrix}
 1 \\
 2
\end{bmatrix}""\begin{bmatrix}
 -2 \\
 3
\end{bmatrix}"


Rank of A= 2

Nullity of A=3


Basis of Null space

From rref of (A)

X1 +17/7X3 -17/7X4 +16/7X5 =0

X2- 9/7X3+ 2/7X4 -6/7X5=0

Let X3, X4 and X5 = 1

X1 -17/7 17/7 -16/7

X2 9/7 -2/7 6/7

X3 = X3 1 + X4 0 + X5 0

X4 0 1 0

X5 0 0 1


Null space has basis formed from

-17/7 17/7 -16/7

9/7 -2/7 6/7

[ 1 , 0 , 0 ]

0 1 0

0 0 1


Augment A by a column e.g "\begin{bmatrix}
 7 \\
 2
\end{bmatrix}"

(A|B)= [ 1 -2 5 -3 4 |7 ]

[ 2 3 1 -4 2 |2 ]


rref(A|B)= [1 0 17/7 -17/7 16/7 |25/7]

[ 0 1 -9/7 2/7 -6/7 |-12/7]

Rank of Augmented A =2

Nullity of Augmented A=4